How many grams of oxygen, O, are in 22.93 g of Al(ClO2)3? Please explain
Answer #1

molecular weight of Al(ClO2)3 = 229.33 gm/mole

oxygen(O) present = 2*3 = 6

molar mass of oxygen= 16 gm/mole

therfore,

in 229.33 gm Al(ClO2)3, Oxygen present = 16*6 = 96 gm

in 22.93 gm Al(ClO2)3, Oxygen present = (96*22.93)/(229.33) = 9.6 gm (answer)